# Tensor network influence functionals in the continuous-time limit: connections to quantum embedding, bath discretization, and higher-order time propagation

### v1

@@ Abstract
We describe two developments of tensor network influence functionals (in particular, influence functional matrix product states (IF-MPS)) for quantum impurity dynamics within the fermionic setting of the Anderson impurity model. The first provides the correct extension of the IF-MPS to continuous time by introducing a related mathematical object, the boundary influence functional MPS. The second connects the dynamics described by a compressed IF-MPS to that of a quantum embedding method with a time-dependent effective bath undergoing non-unitary dynamics. Using these concepts, we implement higher-order time propagators for the quench dynamics of the Anderson impurity model within the boundary IF-MPS formalism. The calculations illustrate the ability of the current formulation to efficiently remove the time-step error in standard discrete-time IF-MPS implementations as well as to interface with state-vector propagation techniques. They also show the advantages of IF-MPS dynamics, with its associated highly compact effective bath dynamics, over state-vector propagation with a static bath discretization.
@@ I Introduction
Quantum impurity models, such as the Anderson impurity model (AIM), consist of an interacting impurity coupled to a (possibly non-interacting) bath. They provide simple settings in which to study quantum many-body physics, and the challenge of describing non-equilibrium dynamics in such models has spurred the development of many computational techniques [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12].
@@
One way to describe quantum impurity dynamics is to use the equation of motion of the impurity reduced density operator obtained by tracing out the bath. The influence of the bath is clearly expressed in a path integral language via the Feynman-Vernon influence functional (IF) [13], which reweights the paths in the path integral. Although the influence functional for a non-interacting bath with linear coupling can be expressed in a compact mathematical form, its effect on the impurity dynamics must still be numerically approximated in practice [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 22, 23].
@@
Recently, tensor network methods have been explored within the IF language to overcome some limitations of other numerical IF techniques, and in particular, to capture longer time memory effects [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. By treating the IF as a temporal wavefunction expressed as a temporal matrix product state (MPS), one can exploit low entanglement in time, leading to a compact representation of certain temporal correlations. Such tensor network-based IF methods have been successfully demonstrated in several contexts, including the spin-boson model [28, 29, 30, 31, 32], hard-core bosons [31], one-dimensional spin chains [26, 27, 33], and more recently, in the context of the AIM [34, 35, 36].
@@
The current work is concerned with two developments of the tensor network IF approach. Although the ideas generalize to IFs of any linearly-coupled non-interacting bath, for concreteness we work specifically with the AIM. The first development removes the time-step error in the standard IF treatment obtained from a Trotterized representation of the path integral. Formulating the problem in the continuous-time limit yields a continuous MPS version of the IF-MPS. We show that the standard IF-MPS in fact does not have a useful continuous-time limit, and instead one must consider a closely related object, the boundary IF-MPS, to define this limit. We provide an explicit construction of the continuous-time boundary IF-MPS for the AIM.
@@
The second development is concerned with the relationship between the (boundary) IF-MPS and discrete bath dynamics. Standard state-vector methods for the dynamics of the AIM use a discrete bath with a fixed set of bath energies and couplings [1, 2, 3, 4, 5]. We show that impurity dynamics defined by an IF-MPS of fixed bond dimension is equivalent to propagating in a space of effective bath orbitals in Liouville space, where the bath energies and couplings dynamically vary with time. This naturally connects tensor network influence functionals to dynamical quantum embedding theories, such as real-time density matrix embedding, which also utilize a dynamical set of bath energies and couplings [37]. An important difference, however, is that the IF-MPS defines a non-unitary dynamics in the bath.
@@
We implement the boundary IF-MPS propagation for the AIM in both the discrete-time and continuous-time formulations. Directly comparing to standard static bath discretizations, we show that the IF-MPS method converges extremely rapidly with respect to the effective bath size, and shows none of the discretization artifacts of standard bath discretizations. Further, the combination of the above two insights suggests that the boundary IF-MPS dynamics in the continuum limit can be efficiently implemented using standard state-vector time-propagation techniques. We use this to implement a high-order Runge-Kutta time-propagator for boundary IF-MPS dynamics and demonstrate the high-order error with time-step. In contrast, we show that higher than first-order Trotter methods in the standard discrete-time IF-MPS still suffer from first-order time-step errors, due to the IF-MPS bond truncation. By Trotterizing the correct continuous-time dynamics, we derive a corrected version of the Trotterized error with the correct time-order scaling, significantly improving on the second-order Trotter formulation currently used in tensor network IF approaches.
@@
The paper is organized as follows. In Sec. II, we first recapitulate the formulation of the IF-MPS for non-interacting fermionic baths, describing in detail the formalism we use here in terms of number-conserving Slater determinants. Through this picture, we establish the connection between the IF-MPS dynamics and Liouville state-vector propagation of an impurity coupled to a set of effective bath orbitals, that is, the dynamics of a quantum embedding of the impurity. In Sec. III, we analyze the continuous-time limit of the IF-MPS and show that the correct object to consider is the boundary IF-MPS and we provide its continuous-time limit. Using the state-vector formalism, we rewrite the continuous-time propagation in terms of a differential equation of motion for the Liouville state-vector in the quantum embedding space. In Sec. IV, we provide numerical results using the boundary IF-MPS for the single impurity Anderson model and compare the discretization errors associated with a static set of bath orbitals with the dynamic set of bath orbitals defined by the IF-MPS. We further analyze the time-step errors from both the discrete-time and continuous-time formulations. In Sec. V, we conclude with discussions of some implications of our results.
@@ II.1 Influence Functional for Single Impurity Anderson Model
where $\hat{d}^{{\dagger}}_{\sigma}$ ($\hat{c}^{{\dagger}}_{i,\sigma}$) creates a fermion of spin $\sigma\in\{\uparrow,\downarrow\}$ in the impurity (bath) orbitals and $\hat{n}_{\sigma}=\hat{d}^{{\dagger}}_{\sigma}\hat{d}_{\sigma}$ is the number density operator of the impurity orbital of spin $\sigma$. Note that $\hat{H}_{SB}$ and $\hat{H}_{B}$ are of noninteracting (quadratic) form. Hereafter, we assume a discrete and finite set of bath orbitals and also assume that the impurity is initially decoupled from the bath, $\hat{\rho}(0)=\hat{\rho}_{S}(0)\otimes\hat{\rho}_{B}$. We further assume a Gaussian initial bath, in particular, the thermal state $\hat{\rho}_{B}\propto e^{-\beta\hat{H}_{B}}$ with inverse temperature $\beta$. For quadratic operators, we omit the hats when expressing their matrix elements in a single-particle basis.
@@
we instead will adopt the super-fermion representation of Liouville space [38, 39, 40, 41, 17]. This uses a super-Fock space with twice the number of orbitals as the original Hilbert space, obtained by applying a particle-hole transformation to the bra of the density operator, and assuming the resulting operator acts on a vacuum to produce a state. We denote states in the super-Fock space using a double bra/ket notation. For example, the initial density operator is written as $\mathinner{|{\rho(0)}\rangle\!\rangle}=\mathinner{|{\rho_{S}(0)}\rangle\!\rangle}\otimes\mathinner{|{\rho_{B}}\rangle\!\rangle}$. The von Neumann equation now becomes a Hamiltonian time evolution with the Liouville operator, $\hat{L}$,
@@
The initial bath state in the super-Fock space, $\mathinner{|{\rho_{B}}\rangle\!\rangle}$, is given by a Slater determinant
@@
and the amplitude of a configuration $s$ in $\rho_{S}$ can be expressed as,
@@
where $t_{N_{t}}=N_{t}\Delta t$ and $\hat{U}^{S}$ ($\hat{U}^{SB}$) is the time evolution operator for $\hat{L}_{S}$ ($\hat{L}_{SB}$ and $\hat{L}_{B}$). The tensor network diagram for this time evolution is shown in Fig. 1, where we see that the $\hat{U}^{S}$ tensor is applied only within the impurity $S$, whereas the tensor for $\hat{U}^{SB}$ is applied to both $S$ and $B$.
Figure 1: A schematic diagram for the real-time evolution of the Anderson Impurity model in Liouville space after $N_{t}=5$ time-steps. The initial density operator is given by $\hat{\rho}(0)=\hat{\rho}_{S}\otimes\hat{\rho}_{B}$, described by a vectorized state in a Liouville space. The time evolution of the density operator is described by time evolution operators following a second-order Trotter decomposition, alternating between $\hat{U}^{S}$ (squares) and $\hat{U}^{SB}$ (rectangles). After the time evolution, the bath degrees of freedom are traced out, which is equivalent to applying the trace tensor, $\text{Tr}_{B}$, to the bath. The influence functional (IF) tensor corresponds to the tensor after the contraction of the bath modes in all $\hat{U}^{SB}$, $\hat{\rho}_{B}$, and $\text{Tr}_{B}$ tensors.
The influence functional (IF) tensor is defined as the tensor arising from the contraction of all the bath degrees of freedom in $\hat{U}_{SB}$, the initial bath density operator, and the trace vector (Fig. 1). This IF tensor is indexed by the state of the impurity orbitals before and after each $\hat{U}_{SB}$, for $N_{t}$ time-steps (denoted $s^{i}_{m}$ and $s^{f}_{m}$, respectively, for the $m$-th time-step) and so is indexed by the configuration of $2N_{t}$ impurity orbitals. In addition, we mention that the IF tensors for different spins are constructed separately thanks to the absence of spin-mixing terms in $\hat{L}_{SB}$, and hence, the spin indices are omitted for brevity.
@@
In the above MPS representation, the bond dimension is given by the dimension of the super-Fock bath space, which, in many cases, is too large to deal with directly. Previous studies [35, 36] have made use of the fermionic Gaussian properties of the IF (arising from the linear coupling and quadratic bath) to find a compressed form of the MPS representation [42, 43]. In this work, we will use a slightly different language to formulate the MPS compression in terms of finding Schmidt vectors in the bath. This algorithm (described in Sec. II.5) is closely related to that in Ref. [42] and improves on the computational scaling in Ref. [43].
II.2 From Influence Functionals to State-Vector Propagation
Figure 2: General matrix product state (MPS) compression scheme for an influence functional MPS (IF-MPS). The bond dimension of the initial uncompressed MPS is given by the dimension of the bath Liouville space, indexed by bath configurations, $b_{m}$ (left). The MPS needs to be converted into canonical form for an optimal truncation. The gauge matrices, $G_{m}$, are inserted in the bath Liouville space to convert the MPS into the canonical form (middle). Gauge matrices transform the original matrix elements of the MPS into left-normalized matrices, denoted as $L_{m}=G_{m}A_{m}G_{m-1}^{-1}$. Afterward, projectors are inserted into the bath subspace that contains the bath states with the largest singular values (right). The projected bath configurations are denoted as $\tilde{b}_{m}$.
In this work, we will often switch between two equivalent pictures: dynamics encoded by a compressed IF-MPS, and a state-vector propagation corresponding to a quantum embedding. To understand this mapping, we first describe the conventional MPS compression scheme (i.e. without using any Gaussian properties of the bath) and explain how it can be interpreted as a type of projected bath dynamics. (We recall that the IF-MPS can, in general, be expressed through Eq. 10 and 11, for arbitrary system-bath quantum dynamics, i.e. even for an interacting bath).
@@ II.2 From Influence Functionals to State-Vector Propagation
where the matrices, $L_{m}^{s^{i}_{m},s^{f}_{m}}$, satisfying Eq. 13, are called left-normalized matrices. The gauge matrices, $G_{m}$, are inserted in the bond space, or the Liouville bath space of the IF-MPS (Fig. 2). The MPS that is composed of the left-normalized matrices is called left-canonical. The left-canonical MPS is then
@@
$\tilde{A}_{m}$ can be viewed as defining a (non-unitary) evolution in the Liouville space of the system and the bath. Alternatively, the MPS gauging and compression procedure can be seen as a pure projected bath dynamics in the Liouville space, $\hat{G}_{m}^{-1}\mathcal{P}_{m}\hat{G}_{m}$, inserted between the system-bath evolution $\hat{U}^{SB}$. The coarse-graining (i.e. projection) of the bath degrees of freedom is referred to as a quantum embedding, and consequently, the truncated IF-MPS dynamics is a quantum embedding scheme with a dynamically evolved bath, similar to (real-time) density matrix embedding theory (DMET) [45, 46, 37, 47], but with the important difference that the dynamics of the projected bath is non-unitary, as $\hat{G}_{m}$ is non-unitary.
As the time-step $\Delta t\to 0$, we are led to the continuous-time limit of the IF-MPS. We consider the subtleties of continuous-time construction in Sec. III.1. However, the above shows that we can also view time evolution as being performed on an embedded state-vector (“wavefunction”) in the projected Liouville space (embedding space) $\{|s_{m}^{i}\tilde{b}_{m}\rangle\rangle\}$. We can then formulate the continuous-time dynamics in terms of the equations of motion for the embedded wavefunction and the bath projectors (or equivalently, the bath density matrix). The dynamics of such a bath density matrix has previously been considered in the MPS language in Ref. [27] for a 1D spin chain, which derived the dissipative contribution of the system-bath coupling to the density matrix dynamics.
We will be interested in the above constructions for the case of a fermionic Gaussian bath where we can replace the discussion of many-body states and density matrices with orbitals and 1-particle reduced density matrices (1-RDM). We now turn to the formulation of the IF-MPS operations in terms of these quantities.

### v2

@@ Abstract
We describe two developments of tensor network influence functionals (in particular, influence functional matrix product states (IF-MPS)) for quantum impurity dynamics within the fermionic setting of the Anderson impurity model. The first provides the correct extension of the IF-MPS to continuous time by introducing a related mathematical object, the boundary influence functional MPS. The second connects the dynamics described by a compressed IF-MPS to that of a quantum embedding method with a time-dependent effective bath undergoing nonunitary dynamics. Using these concepts, we implement higher-order time propagators for the quench dynamics of the Anderson impurity model within the boundary IF-MPS formalism. The calculations illustrate the ability of the current formulation to efficiently remove the time step error in standard discrete-time IF-MPS implementations as well as to interface with state vector propagation techniques. They also show the advantages of IF-MPS dynamics, with its associated highly compact effective bath dynamics, over state vector propagation with a static bath discretization.
@@ I Introduction
Quantum impurity models, such as the Anderson impurity model (AIM), consist of an interacting impurity coupled to a (possibly noninteracting) bath. They provide simple settings in which to study quantum many-body physics, including the Kondo effect in magnetic impurities [1], and serve as a starting point of computational embedding frameworks, such as dynamical mean-field theory [2] or density matrix embedding theory [3, 4], to solve lattice problems. The challenge of describing non-equilibrium dynamics in such models has spurred the development of many computational techniques [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16].
@@
One way to describe quantum impurity dynamics is to use the equation of motion of the impurity reduced density operator obtained by tracing out the bath. The influence of the bath is clearly expressed in a path integral language via the Feynman-Vernon influence functional (IF) [17], which reweights the paths in the path integral. Although the influence functional for a noninteracting bath with linear coupling can be expressed in a compact mathematical form, its effect on the impurity dynamics must still be numerically approximated in practice [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 26, 27].
@@
Recently, tensor network methods have been explored within the IF language to overcome some limitations of other numerical IF techniques, and in particular, to capture longer-time memory effects [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42]. By treating the IF as a temporal wavefunction expressed as a temporal matrix product state (MPS), one can exploit low entanglement in time, leading to a compact representation of certain temporal correlations. Such tensor network-based IF methods have been successfully demonstrated in several contexts, including the spin-boson model [32, 33, 34, 35, 36], hard-core bosons [35], one-dimensional spin chains [30, 31, 37], and more recently, in the context of the AIM [38, 39, 40, 41, 42, 43].
@@
The current paper is concerned with two developments of the tensor network IF approach. Although the ideas generalize to IFs of any linearly coupled noninteracting bath, for concreteness we work specifically with the AIM, which expands on the previous developments based on the standard IF-MPS framework [38, 39, 40, 41]. The first development removes the time step error in the standard IF treatment obtained from a Trotterized representation of the path integral. Formulating the problem in the continuous-time limit yields a continuous MPS version of the IF-MPS. We show that the standard IF-MPS in fact does not have a useful continuous-time limit, and instead one must consider a closely related object, the boundary IF-MPS, to define this limit. We provide an explicit construction of the continuous-time boundary IF-MPS for the AIM.
@@
The second development is concerned with the relationship between the (boundary) IF-MPS and discrete bath dynamics. Standard state vector methods for the dynamics of the AIM use a discrete bath with a fixed set of bath energies and couplings [5, 6, 7, 8, 9]. We show that impurity dynamics defined by an IF-MPS of fixed bond dimension is equivalent to propagating in a space of effective bath orbitals in Liouville space, where the bath energies and couplings dynamically vary with time. This naturally connects tensor network influence functionals to dynamical quantum embedding theories, such as real-time density matrix embedding, which also utilize a dynamical set of bath energies and couplings [44]. An important difference, however, is that the IF-MPS defines a nonunitary dynamics in the bath.
@@
We implement the boundary IF-MPS propagation for the AIM in both the discrete-time and continuous-time formulations. Directly comparing to standard static bath discretizations, we show that the IF-MPS method converges extremely rapidly with respect to the effective bath size, and shows none of the discretization artifacts of standard bath discretizations. Further, the combination of the above two insights suggests that the boundary IF-MPS dynamics in the continuum limit can be efficiently implemented using standard state vector time-propagation techniques. We use this to implement a high-order Runge-Kutta time propagator for boundary IF-MPS dynamics and demonstrate the high-order error with time step. In contrast, we show that higher than first-order Trotter methods in the standard discrete-time IF-MPS still suffer from first-order time step errors, due to the IF-MPS bond truncation. By Trotterizing the correct continuous-time dynamics, we derive a corrected version of the Trotterized error with the correct time-order scaling, significantly improving on the second-order Trotter formulation currently used in tensor network IF approaches.
@@
The paper is organized as follows. In Sec. II, we first recapitulate the formulation of the IF-MPS for noninteracting fermionic baths, describing in detail the formalism we use here in terms of number-conserving Slater determinants. Through this picture, we establish the connection between the IF-MPS dynamics and Liouville state vector propagation of an impurity coupled to a set of effective bath orbitals, that is, the dynamics of a quantum embedding of the impurity. In Sec. III, we analyze the continuous-time limit of the IF-MPS and show that the correct object to consider is the boundary IF-MPS and we provide its continuous-time limit. Using the state vector formalism, we rewrite the continuous-time propagation in terms of a differential equation of motion for the Liouville state vector in the quantum embedding space. In Sec. IV, we provide numerical results using the boundary IF-MPS for the single impurity Anderson model and compare the discretization errors associated with a static set of bath orbitals with the dynamic set of bath orbitals defined by the IF-MPS. We further analyze the time step errors from both the discrete-time and continuous-time formulations. In Sec. V, we conclude with discussions of some implications of our results.
@@ II.1 Influence functional for single impurity Anderson model
[…]
@@
we instead will adopt the super-fermion representation of Liouville space [46, 47, 48, 49, 21]. This uses a super-Fock space with twice the number of orbitals as the original Hilbert space, obtained by applying a particle-hole transformation to the bra of the density operator, and assuming the resulting operator acts on a vacuum to produce a state. We denote states in the super-Fock space using a double bra/ket notation. For example, the initial density operator is written as $\mathinner{|{\rho(0)}\rangle\!\rangle}=\mathinner{|{\rho_{S}(0)}\rangle\!\rangle}\otimes\mathinner{|{\rho_{B}}\rangle\!\rangle}$. The von Neumann equation now becomes a Hamiltonian time evolution with the Liouville operator $\hat{L}$,
@@
The initial bath state in the super-Fock space $\mathinner{|{\rho_{B}}\rangle\!\rangle}$ is given by a Slater determinant
@@
and the amplitude of a configuration $s$ in $\rho_{S}$ can be expressed as
@@
where $t_{N_{t}}=N_{t}\Delta t$ and $\hat{U}^{S}$ ($\hat{U}^{SB}$) is the time-evolution operator for $\hat{L}_{S}$ ($\hat{L}_{SB}$ and $\hat{L}_{B}$). The tensor network diagram for this time evolution is shown in Fig. 1, where we see that the $\hat{U}^{S}$ tensor is applied only within the impurity $S$, whereas the tensor for $\hat{U}^{SB}$ is applied to both $S$ and $B$.
Figure 1: A schematic diagram for the real-time evolution of the Anderson impurity model in Liouville space after $N_{t}=5$ time steps. The initial density operator is given by $\hat{\rho}(0)=\hat{\rho}_{S}\otimes\hat{\rho}_{B}$, described by a vectorized state in a Liouville space. The time evolution of the density operator is described by time-evolution operators following a second-order Trotter decomposition, alternating between $\hat{U}^{S}$ (squares) and $\hat{U}^{SB}$ (rectangles). After the time evolution, the bath degrees of freedom are traced out, which is equivalent to applying the trace tensor $\text{Tr}_{B}$ to the bath. The influence functional (IF) tensor corresponds to the tensor after the contraction of the bath modes in all $\hat{U}^{SB}$, $\hat{\rho}_{B}$, and $\text{Tr}_{B}$ tensors, grouped with the red-dashed lines.
The influence functional (IF) tensor is defined as the tensor arising from the contraction of all the bath degrees of freedom in $\hat{U}_{SB}$, the initial bath density operator, and the trace vector (Fig. 1). This IF tensor is indexed by the state of the impurity orbitals before and after each $\hat{U}_{SB}$, for $N_{t}$ time steps (denoted $s^{i}_{m}$ and $s^{f}_{m}$, respectively, for the $m$-th time step) and so is indexed by the configuration of $2N_{t}$ impurity orbitals. In addition, we mention that the IF tensors for different spins are constructed separately thanks to the absence of spin-mixing terms in $\hat{L}_{SB}$, and hence, the spin indices are omitted for brevity.
@@
In the above MPS representation, the bond dimension is given by the dimension of the super-Fock bath space, which, in many cases, is too large to deal with directly. Previous studies [39, 40] have made use of the fermionic Gaussian properties of the IF (arising from the linear coupling and quadratic bath) to find a compressed form of the MPS representation [50, 51]. In this paper, we will use a slightly different language to formulate the MPS compression in terms of finding Schmidt vectors in the bath. This algorithm (described in Sec. II.5) is closely related to that in Ref. [50] and improves on the computational scaling in Ref. [51].
II.2 From influence functionals to state vector propagation
Figure 2: General matrix product state (MPS) compression scheme for an influence functional MPS (IF-MPS). The bond dimension of the initial uncompressed MPS is given by the dimension of the bath Liouville space, indexed by bath configurations, $b_{m}$ (left). The MPS needs to be converted into canonical form for an optimal truncation. The gauge matrices $G_{m}$ are inserted in the bath Liouville space to convert the MPS into the canonical form (middle). Gauge matrices transform the original matrix elements of the MPS into left-normalized matrices, denoted as $L_{m}=G_{m}A_{m}G_{m-1}^{-1}$. Afterward, projectors are inserted into the bath subspace that contains the bath states with the largest singular values (right). The projected bath configurations are denoted as $\tilde{b}_{m}$.
In this paper, we will often switch between two equivalent pictures: dynamics encoded by a compressed IF-MPS, and a state vector propagation corresponding to a quantum embedding. To understand this mapping, we first describe the conventional MPS compression scheme (i.e., without using any Gaussian properties of the bath) and explain how it can be interpreted as a type of projected bath dynamics. (We recall that the IF-MPS can, in general, be expressed through Eqs. 10 and 11, for arbitrary system-bath quantum dynamics, i.e., even for an interacting bath).
@@ II.2 From influence functionals to state vector propagation
where the matrices $L_{m}^{s^{i}_{m},s^{f}_{m}}$ satisfying Eq. 13, are called left-normalized matrices. The gauge matrices, $G_{m}$, are inserted in the bond space, or the Liouville bath space of the IF-MPS (Fig. 2). The MPS that is composed of the left-normalized matrices is called left-canonical. The left-canonical MPS is then
@@
$\tilde{A}_{m}$ can be viewed as defining a (nonunitary) evolution in the Liouville space of the system and the bath. Alternatively, the MPS gauging and compression procedure can be seen as a pure projected bath dynamics in the Liouville space $\hat{G}_{m}^{-1}\mathcal{P}_{m}\hat{G}_{m}$, inserted between the system-bath evolution $\hat{U}^{SB}$. The coarse graining (i.e., projection) of the bath degrees of freedom is referred to as a quantum embedding, and consequently, the truncated IF-MPS dynamics is a quantum embedding scheme with a dynamically evolved bath, similar to (real-time) density matrix embedding theory (DMET) [3, 4, 44, 53], but with the important difference that the dynamics of the projected bath is nonunitary, as $\hat{G}_{m}$ is nonunitary.
As the time step $\Delta t\to 0$, we are led to the continuous-time limit of the IF-MPS. We consider the subtleties of continuous-time construction in Sec. III.1. However, the above shows that we can also view time evolution as being performed on an embedded state vector (“wavefunction”) in the projected Liouville space (embedding space) $\{|s_{m}^{i}\tilde{b}_{m}\rangle\rangle\}$. We can then formulate the continuous-time dynamics in terms of the equations of motion for the embedded wavefunction and the bath projectors (or equivalently, the bath density matrix). The dynamics of such a bath density matrix has previously been considered in the MPS language in Ref. [31] for a 1D spin chain, which derived the dissipative contribution of the system-bath coupling to the density matrix dynamics.
We will be interested in the above constructions for the case of a fermionic Gaussian bath where we can replace the discussion of many-body states and density matrices with orbitals and one-particle reduced density matrices (1-RDM). We now turn to the formulation of the IF-MPS operations in terms of these quantities.